# atomic orbital API Application Programmer Interface

### v1

@@ 4.1. Algorithm
$\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}\right)\delta\varphi_{2}+\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}\right)\varphi_{2}\\ +c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}+c_{2}^{2}\left(V_{\text{nuc}}+J(22)\right)\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{1}-\varepsilon_{12}\varphi_{1}+\left(\varepsilon_{22}\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}-\delta\varepsilon_{22}\right)\varphi_{2}\\ +\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(22)\right)\varphi_{2}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{1}=0$ (4.18)
Finally,
$\delta\varphi_{1}=-\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\varphi_{1}-R_{1}\mathfrak{F}_{1}$ (4.19)
$\delta\varphi_{2}=-\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\varphi_{2}-R_{2}\mathfrak{F}_{2}$
@@
$\mathfrak{F}_{2}=-\varepsilon_{12}\delta\varphi_{1}-(\delta\varepsilon_{12}+\varepsilon_{12})\varphi_{1}+\left(\varepsilon_{22}\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}-\delta\varepsilon_{22}\right)\varphi_{2}\\ +\left(V_{\text{nuc}}+J(22)\right)\left(c_{2}^{2}\delta\varphi_{2}+\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\varphi_{2}\right)+J(12)(c_{1}c_{2}\delta\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)\varphi_{1})\\ +2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(1\delta 2)\varphi_{1}$
@@
Overall, the complete Newton procedure can be schematically described by Figure 3. The two orbital MCSCF for the helium atom results in the coefficients $c_{1}=0.99793,c_{2}=-0.06430$, associated to $\theta=-0.06435$, and orbitals symmetric with respect to origin, shown in Figure 1. The corresponding total energy is $-2.87799$.
Figure 1. Two determinants ground state of the helium atom with the coefficients $c_{0}=0.99793$ and $c_{1}=-0.06430$, correspondingly. The corresponding total energy is $-2.87799$.
@@ 5. Symmetric hessian and shift
$\frac{\delta\mathcal{L}}{\delta\varphi_{1}}=c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}$ (5.1)
$\frac{\delta\mathcal{L}}{\delta\varphi_{2}}=c_{2}^{2}\left(h+\frac{1}{|x|}*\varphi_{2}^{2}\right)\varphi_{2}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{1}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2}$
@@
$\frac{\partial\mathcal{L}}{\partial\varepsilon_{11}}$ $=-\frac{1}{2}\left(\left\|\varphi_{1}\right\|^{2}-1\right)$ (5.2)
$\frac{\partial\mathcal{L}}{\partial\varepsilon_{22}}$ $=-\frac{1}{2}\left(\left\|\varphi_{2}\right\|^{2}-1\right)$
$\frac{\partial\mathcal{L}}{\partial\varepsilon_{12}}$ $=-\int\varphi_{1}\varphi_{2}$
@@
with respect $\delta w\in W$. This is a linear equation that we solve iteratively and accelerating with DIIS. The Jacobian $d\mathcal{R}(w)$ equals to the Hessian $d\nabla\mathcal{L}(w)$. Then the update is $w_{n+1}=w_{n}+\delta w$.
$d\mathcal{R}_{i}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{i}(w)\delta\varphi_{1}+\ldots+\frac{\partial\mathcal{R}_{i}}{\partial\varepsilon_{12}}(w)\delta\varepsilon_{12},\quad i=1,2,$
and the rest
Similarly, $d\mathcal{R}_{2}(w)\delta w$ consists of
and the remaining terms
$\frac{\partial\mathcal{R}_{2}}{\partial\varepsilon_{11}}(w)\delta\varepsilon_{11}+\frac{\partial\mathcal{R}_{2}}{\partial\varepsilon_{22}}(w)\delta\varepsilon_{22}+\frac{\partial\mathcal{R}_{2}}{\partial\varepsilon_{12}}(w)\delta\varepsilon_{12}=-\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}$
@@
$\partial_{\varphi_{1}}\mathcal{R}_{3}(w)\delta\varphi_{1}=-\sin 2\theta((\delta 1|h|1)+(\delta 11|11))+\cos 2\theta(\delta 12|12)$
@@
$\frac{\partial\mathcal{R}_{3}}{\partial\theta}(w)\delta\theta=\left(-\frac{1}{2}\cos 2\theta(2(1|h|1)+(11|11))+\frac{1}{2}\cos 2\theta(2(2|h|2)+(22|22))-\sin 2\theta(12|12)\right)\delta\theta$
@@
$d\mathcal{R}_{4}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{4}(w)\delta\varphi_{1}=-\int\varphi_{1}\delta\varphi_{1}$
$d\mathcal{R}_{5}(w)\delta w=\partial_{\varphi_{2}}\mathcal{R}_{5}(w)\delta\varphi_{2}=-\int\varphi_{2}\delta\varphi_{2}$
$d\mathcal{R}_{6}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{6}(w)\delta\varphi_{1}+\partial_{\varphi_{2}}\mathcal{R}_{6}(w)\delta\varphi_{2}=-\int\delta\varphi_{1}\varphi_{2}-\int\varphi_{1}\delta\varphi_{2}$
@@ 5.1. Self-consistent form
$\delta\theta=-\left(\frac{\partial\mathcal{R}_{3}}{\partial\theta}(w)+\lambda\right)^{-1}\left(\partial_{\varphi_{1}}\mathcal{R}_{3}(w)\delta\varphi_{1}+\partial_{\varphi_{2}}\mathcal{R}_{3}(w)\delta\varphi_{2}+\mathcal{R}_{3}(w)\right)$
@@
$\delta\varepsilon_{11}\varphi_{1}+\delta\varepsilon_{12}\varphi_{2}+(\varepsilon_{11}-\lambda)\delta\varphi_{1}+\varepsilon_{12}\delta\varphi_{2}=F(\delta\varphi_{1},\delta\varphi_{2},\delta\theta,w)$ (5.3)
$\delta\varepsilon_{12}\varphi_{1}+\delta\varepsilon_{22}\varphi_{2}+\varepsilon_{12}\delta\varphi_{1}+(\varepsilon_{22}-\lambda)\delta\varphi_{2}=G(\delta\varphi_{1},\delta\varphi_{2},\delta\theta,w)$
@@
$\begin{pmatrix}\left\|\varphi_{1}\right\|^{2}+(\varepsilon_{11}-\lambda)\lambda&0&\int\varphi_{1}\varphi_{2}&0&\varepsilon_{12}\\ 0&\left\|\varphi_{2}\right\|^{2}+(\varepsilon_{22}-\lambda)\lambda&\int\varphi_{1}\varphi_{2}&\varepsilon_{12}&0\\ \int\varphi_{1}\varphi_{2}&0&\left\|\varphi_{2}\right\|^{2}&\varepsilon_{11}-\lambda&0\\ 0&\int\varphi_{1}\varphi_{2}&\left\|\varphi_{1}\right\|^{2}&0&\varepsilon_{22}-\lambda\\ 0&0&-\lambda&1&1\end{pmatrix}\begin{pmatrix}\delta\varepsilon_{11}\\ \delta\varepsilon_{22}\\ \delta\varepsilon_{12}\\ \int\delta\varphi_{1}\varphi_{2}\\ \int\varphi_{1}\delta\varphi_{2}\end{pmatrix}\\ =\begin{pmatrix}\int F\varphi_{1}-\frac{1}{2}(\varepsilon_{11}-\lambda)\left(1-\left\|\varphi_{1}\right\|^{2}\right)\\ \int G\varphi_{2}-\frac{1}{2}(\varepsilon_{22}-\lambda)\left(1-\left\|\varphi_{2}\right\|^{2}\right)\\ \int F\varphi_{2}-\frac{1}{2}\varepsilon_{12}\left(1-\left\|\varphi_{2}\right\|^{2}\right)\\ \int G\varphi_{1}-\frac{1}{2}\varepsilon_{12}\left(1-\left\|\varphi_{1}\right\|^{2}\right)\\ -\int\varphi_{1}\varphi_{2}\end{pmatrix}$ (5.4)
We search for orthonormal orbitals $\varphi_{1},\varphi_{2}$. Therefore, the determinant of this system should stay close to $-\varepsilon_{11}-\varepsilon_{22}$ during each Newton’s update, provided $\lambda=0$. In particular, the matrix should be invertible for each Newton’s iteration $w=w_{n}$.
$\int F\varphi_{1}=c_{1}^{2}(\delta 1|h|1)+3c_{1}^{2}(\delta 11|11)+c_{1}c_{2}(\delta 12|12)+2c_{1}c_{2}(1\delta 2|12)\\ +(c_{1}^{2}-\delta\theta\sin 2\theta)((1|h|1)+(11|11))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(12|12)-\varepsilon_{11}\left\|\varphi_{1}\right\|^{2}-\varepsilon_{12}\int\varphi_{1}\varphi_{2}$ (5.5)
$\int F\varphi_{2}=c_{1}^{2}(\delta 1|h|2)+c_{1}^{2}(\delta 12|11)+2c_{1}^{2}(\delta 11|12)+c_{1}c_{2}(\delta 12|22)+c_{1}c_{2}(1\delta 2|22)+c_{1}c_{2}(2\delta 2|12)\\ +(c_{1}^{2}-\delta\theta\sin 2\theta)((1|h|2)+(11|12))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(12|22)-\varepsilon_{11}\int\varphi_{1}\varphi_{2}-\varepsilon_{12}\left\|\varphi_{2}\right\|^{2}$ (5.6)
$\int G\varphi_{2}=c_{2}^{2}(\delta 2|h|2)+3c_{2}^{2}(\delta 22|22)+2c_{1}c_{2}(\delta 12|12)+c_{1}c_{2}(1\delta 2|12)\\ +(c_{2}^{2}+\delta\theta\sin 2\theta)((2|h|2)+(22|22))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(12|12)-\varepsilon_{22}\left\|\varphi_{2}\right\|^{2}-\varepsilon_{12}\int\varphi_{1}\varphi_{2}$ (5.7)
$\int G\varphi_{1}=c_{2}^{2}(\delta 2|h|1)+c_{2}^{2}(1\delta 2|22)+2c_{2}^{2}(\delta 22|12)+c_{1}c_{2}(\delta 12|11)+c_{1}c_{2}(1\delta 2|11)+c_{1}c_{2}(1\delta 1|12)\\ +(c_{2}^{2}+\delta\theta\sin 2\theta)((1|h|2)+(12|22))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(11|12)-\varepsilon_{22}\int\varphi_{1}\varphi_{2}-\varepsilon_{12}\left\|\varphi_{1}\right\|^{2}$ (5.8)
@@
$\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}+\lambda\right)\delta\varphi_{1}+\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}+\lambda\right)\varphi_{1}\\ +c_{1}^{2}\left(V_{\text{nuc}}+J(11)\right)\delta\varphi_{1}+2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{2}-\varepsilon_{12}\varphi_{2}-\left((\varepsilon_{11}-\lambda)\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}+\delta\varepsilon_{11}+\lambda\right)\varphi_{1}\\ +\left(c_{1}^{2}-\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(11)\right)\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{2}=0$ (5.9)
$\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}+\lambda\right)\delta\varphi_{2}+\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}+\lambda\right)\varphi_{2}\\ +c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}+c_{2}^{2}\left(V_{\text{nuc}}+J(22)\right)\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{1}-\varepsilon_{12}\varphi_{1}+\left((\varepsilon_{22}-\lambda)\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}-\delta\varepsilon_{22}-\lambda\right)\varphi_{2}\\ +\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(22)\right)\varphi_{2}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{1}=0$ (5.10)
Finally, introducing
@@
$\delta\varphi_{1}=-\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\varphi_{1}-R_{1}^{\lambda}\mathfrak{F}_{1}^{\lambda}$ (5.11)
$\delta\varphi_{2}=-\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\varphi_{2}-R_{2}^{\lambda}\mathfrak{F}_{2}^{\lambda}$
@@
$\mathfrak{F}_{2}^{\lambda}=-\varepsilon_{12}\delta\varphi_{1}-(\delta\varepsilon_{12}+\varepsilon_{12})\varphi_{1}+\left((\varepsilon_{22}-\lambda)\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}-\delta\varepsilon_{22}-\lambda\right)\varphi_{2}\\ +\left(V_{\text{nuc}}+J(22)\right)\left(c_{2}^{2}\delta\varphi_{2}+\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\varphi_{2}\right)+J(12)(c_{1}c_{2}\delta\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)\varphi_{1})\\ +2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(1\delta 2)\varphi_{1}$

### v2

@@ 3.1. Self-consistent form
(3.19) $\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}\right)\delta\varphi_{2}+\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}\right)\varphi_{2}\\ +c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}+c_{2}^{2}\left(V_{\text{nuc}}+J(22)\right)\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{1}-\varepsilon_{12}\varphi_{1}+\left(\varepsilon_{22}\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}-\delta\varepsilon_{22}\right)\varphi_{2}\\ +\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(22)\right)\varphi_{2}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{1}=0.$
Finally, inverting the kinetic energy we obtain the system
(3.20) $\delta\varphi_{1}=-\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\varphi_{1}-R_{1}\mathfrak{F}_{1},$
$\delta\varphi_{2}=-\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\varphi_{2}-R_{2}\mathfrak{F}_{2},$
where
@@
and
$\mathfrak{F}_{2}=-\varepsilon_{12}\delta\varphi_{1}-(\delta\varepsilon_{12}+\varepsilon_{12})\varphi_{1}+\left(\varepsilon_{22}\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}-\delta\varepsilon_{22}\right)\varphi_{2}\\ +\left(V_{\text{nuc}}+J(22)\right)\left(c_{2}^{2}\delta\varphi_{2}+\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\varphi_{2}\right)+J(12)(c_{1}c_{2}\delta\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)\varphi_{1})\\ +2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(1\delta 2)\varphi_{1}.$
This completes the description of the self-consistent form (2.4), suitable for numerical treatment of the Newton equation (2.3) in the two-determinant case. These equations can be easily discretized using multiwavelets.
[…]
Figure 1. Newton energy optimization step.
Figure 2. Two-determinant ground state of the helium atom with the coefficients $c_{0}=0.99793$ and $c_{1}=-0.06430$, correspondingly. The corresponding total energy is 2.87799-2.87799.
@@
Overall, the complete Newton procedure can be schematically described by Figure 1. It is not necessary to perform the Löwdin transformation and the optimization of CI coefficients after every Newton iteration; however, including these steps significantly accelerates convergence. In other words, the dominant factor affecting convergence is the treatment of normalization. To achieve second-order convergence, the orbitals must be normalized after each outer-loop iteration. The two orbital MCSCF for the helium atom results in the coefficients $c_{1}=0.99793,c_{2}=-0.06430$, associated to $\theta=-0.06435$, and orbitals symmetric with respect to origin, shown in Figure 2. The corresponding total energy is 2.87799-2.87799.
@@ 4. Symmetric hessian and shift
(4.1) $\frac{\delta\mathcal{L}}{\delta\varphi_{1}}=c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2},$
$\frac{\delta\mathcal{L}}{\delta\varphi_{2}}=c_{2}^{2}\left(h+\frac{1}{|x|}*\varphi_{2}^{2}\right)\varphi_{2}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{1}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2},$
@@
(4.2) $\frac{\partial\mathcal{L}}{\partial\varepsilon_{11}}$ $=-\frac{1}{2}\left(\left\|\varphi_{1}\right\|^{2}-1\right),$
$\frac{\partial\mathcal{L}}{\partial\varepsilon_{22}}$ $=-\frac{1}{2}\left(\left\|\varphi_{2}\right\|^{2}-1\right),$
$\frac{\partial\mathcal{L}}{\partial\varepsilon_{12}}$ $=-\int\varphi_{1}\varphi_{2}.$
@@
with respect $\delta w\in W$. This is a linear equation that we solve iteratively and accelerating with DIIS. The Jacobian $d\mathcal{R}(w)$ equals to the Hessian $d\nabla\mathcal{L}(w)$. As above, the solution of this equation allows to update the current iterate $w_{n}$ by $w_{n+1}=w_{n}+\delta w$. The derivatives of the first two components (3.3) were found in the previous section, and they are exactly the same here for the new $\mathcal{R}$.
@@
$\partial_{\varphi_{1}}\mathcal{R}_{3}(w)\delta\varphi_{1}=-\sin 2\theta((\delta 1|h|1)+(\delta 11|11))+\cos 2\theta(\delta 12|12),$
@@
and
$\frac{\partial\mathcal{R}_{3}}{\partial\theta}(w)\delta\theta=\left(-\frac{1}{2}\cos 2\theta(2(1|h|1)+(11|11))+\frac{1}{2}\cos 2\theta(2(2|h|2)+(22|22))-\sin 2\theta(12|12)\right)\delta\theta.$
@@
$d\mathcal{R}_{4}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{4}(w)\delta\varphi_{1}=-\int\varphi_{1}\delta\varphi_{1},$
$d\mathcal{R}_{5}(w)\delta w=\partial_{\varphi_{2}}\mathcal{R}_{5}(w)\delta\varphi_{2}=-\int\varphi_{2}\delta\varphi_{2},$
$d\mathcal{R}_{6}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{6}(w)\delta\varphi_{1}+\partial_{\varphi_{2}}\mathcal{R}_{6}(w)\delta\varphi_{2}=-\int\delta\varphi_{1}\varphi_{2}-\int\varphi_{1}\delta\varphi_{2}.$
@@ 4.1. Self-consistent form
$\delta\theta=-\left(\frac{\partial\mathcal{R}_{3}}{\partial\theta}(w)+\lambda\right)^{-1}\left(\partial_{\varphi_{1}}\mathcal{R}_{3}(w)\delta\varphi_{1}+\partial_{\varphi_{2}}\mathcal{R}_{3}(w)\delta\varphi_{2}+\mathcal{R}_{3}(w)\right).$
@@
(4.3) $\delta\varepsilon_{11}\varphi_{1}+\delta\varepsilon_{12}\varphi_{2}+(\varepsilon_{11}-\lambda)\delta\varphi_{1}+\varepsilon_{12}\delta\varphi_{2}=F(\delta\varphi_{1},\delta\varphi_{2},\delta\theta,w),$
$\delta\varepsilon_{12}\varphi_{1}+\delta\varepsilon_{22}\varphi_{2}+\varepsilon_{12}\delta\varphi_{1}+(\varepsilon_{22}-\lambda)\delta\varphi_{2}=G(\delta\varphi_{1},\delta\varphi_{2},\delta\theta,w).$
@@
(4.4) $\begin{pmatrix}\left\|\varphi_{1}\right\|^{2}+(\varepsilon_{11}-\lambda)\lambda&0&\int\varphi_{1}\varphi_{2}&0&\varepsilon_{12}\\ 0&\left\|\varphi_{2}\right\|^{2}+(\varepsilon_{22}-\lambda)\lambda&\int\varphi_{1}\varphi_{2}&\varepsilon_{12}&0\\ \int\varphi_{1}\varphi_{2}&0&\left\|\varphi_{2}\right\|^{2}&\varepsilon_{11}-\lambda&0\\ 0&\int\varphi_{1}\varphi_{2}&\left\|\varphi_{1}\right\|^{2}&0&\varepsilon_{22}-\lambda\\ 0&0&-\lambda&1&1\end{pmatrix}\begin{pmatrix}\delta\varepsilon_{11}\\ \delta\varepsilon_{22}\\ \delta\varepsilon_{12}\\ \int\delta\varphi_{1}\varphi_{2}\\ \int\varphi_{1}\delta\varphi_{2}\end{pmatrix}\\ =\begin{pmatrix}\int F\varphi_{1}-\frac{1}{2}(\varepsilon_{11}-\lambda)\left(1-\left\|\varphi_{1}\right\|^{2}\right)\\ \int G\varphi_{2}-\frac{1}{2}(\varepsilon_{22}-\lambda)\left(1-\left\|\varphi_{2}\right\|^{2}\right)\\ \int F\varphi_{2}-\frac{1}{2}\varepsilon_{12}\left(1-\left\|\varphi_{2}\right\|^{2}\right)\\ \int G\varphi_{1}-\frac{1}{2}\varepsilon_{12}\left(1-\left\|\varphi_{1}\right\|^{2}\right)\\ -\int\varphi_{1}\varphi_{2}\end{pmatrix}$
We search for orthonormal orbitals $\varphi_{1},\varphi_{2}$. Therefore, the determinant of this system should stay close to $\varepsilon_{11}-\varepsilon_{22}$ during each Newton’s update, provided $\lambda=0$. In particular, the matrix should be invertible for each Newton’s iteration $w=w_{n}$. The integrals in the right hand side of (4.4) do not depend on $\lambda$ and they are computed by Formulas (3.6)-(3.9).
@@
(4.5) $\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}+\lambda\right)\delta\varphi_{1}+\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}+\lambda\right)\varphi_{1}\\ +c_{1}^{2}\left(V_{\text{nuc}}+J(11)\right)\delta\varphi_{1}+2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{2}-\varepsilon_{12}\varphi_{2}-\left((\varepsilon_{11}-\lambda)\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}+\delta\varepsilon_{11}+\lambda\right)\varphi_{1}\\ +\left(c_{1}^{2}-\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(11)\right)\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{2}=0$
and
(4.6) $\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}+\lambda\right)\delta\varphi_{2}+\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}+\lambda\right)\varphi_{2}\\ +c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}+c_{2}^{2}\left(V_{\text{nuc}}+J(22)\right)\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{1}-\varepsilon_{12}\varphi_{1}+\left((\varepsilon_{22}-\lambda)\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}-\delta\varepsilon_{22}-\lambda\right)\varphi_{2}\\ +\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(22)\right)\varphi_{2}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{1}=0.$
Finally, introducing the resolvents at shifted orbital energies
@@
we can rewrite Equations (4.5), (4.6) as
(4.7) $\delta\varphi_{1}=-\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\varphi_{1}-R_{1}^{\lambda}\mathfrak{F}_{1}^{\lambda},$
$\delta\varphi_{2}=-\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\varphi_{2}-R_{2}^{\lambda}\mathfrak{F}_{2}^{\lambda},$
where
@@
and
$\mathfrak{F}_{2}^{\lambda}=-\varepsilon_{12}\delta\varphi_{1}-(\delta\varepsilon_{12}+\varepsilon_{12})\varphi_{1}+\left((\varepsilon_{22}-\lambda)\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}-\delta\varepsilon_{22}-\lambda\right)\varphi_{2}\\ +\left(V_{\text{nuc}}+J(22)\right)\left(c_{2}^{2}\delta\varphi_{2}+\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\varphi_{2}\right)+J(12)(c_{1}c_{2}\delta\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)\varphi_{1})\\ +2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(1\delta 2)\varphi_{1}.$